A bat strikes a 0.145-kg baseball. Just before impact, the ball is traveling horizontally to the right at 50.0 m/s , and it leaves the bat traveling to the left at an angle of 40 ∘ above horizontal with a speed of 55.0 m/s . The ball and bat are in contact for 1.75 ms .

A) Find the horizontal component of the average force on the ball. Take the x-direction to be positive to the right

B)Find the vertical component of the average force on the ball.
Answer #1

v1 is initial velocity horizontal
v2h is final velocity horizontal
at is acceleration-time velocity change

v1-at =-v2h
(v1+v2h)/t =a
50 +55cos40/.00175 =a
a= -52647.11 m/s/s
F=ma
F = (0.145)(-52647.11)
Fhorizontal average = 7633.83 N to the left

v1=0
v2v =55sin40
v=at
a=v/t
F=ma
F=mv/t
F = (0.145)(55sin40)/(0.00175)
Fvertical average = 3222 N upward